OpenStudy (anonymous):

When you add square roots, do you add the insides? ex. square root of 3 plus the square root of 3 i kinda for got this over the summer XD

OpenStudy (anonymous):

\[\sqrt{3}\] + \[\sqrt{3}\]

OpenStudy (anonymous):

No you cannot.

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

\[\sqrt{3}+\sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

ohh... i got it now

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

and another one: do you reduce and what if the discriminant isnt the same #?

OpenStudy (anonymous):

do you reduce what?